OpenStudy (anonymous):

Consider the planes given by the equations x+y+4y=2, and z-(x+2y)=5. Find a vector V parallel to the line of intersection of the planes.

OpenStudy (anonymous):

x+y+4z=2 and z-(x+2y)=5 right?

OpenStudy (anonymous):

The first think you want to do is get the line of intersection.

OpenStudy (anonymous):

Actually hold on...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay so that's my guess is to find the line of intersection.

OpenStudy (anonymous):

I would do that by solving for z in one equation, and substituting that expression into the other equation.

OpenStudy (anonymous):

okay so z=0.5-1x/4-1y/4

OpenStudy (anonymous):

I would have done z = 5 + x + 2y....

OpenStudy (anonymous):

So we get x+y+4(5+x+2y) = 2

OpenStudy (anonymous):

okay now what?

OpenStudy (anonymous):

Solve for y to get it into y = mx + b form.

OpenStudy (anonymous):

our vector should just be <1, m> I'm thinking.

OpenStudy (anonymous):

We did this the ghetto way though, I have a feeling they wanted you to do it differently.

OpenStudy (anonymous):

We'll see. haha

OpenStudy (anonymous):

What did you end up getting?

OpenStudy (anonymous):

y=-2x-2 Does that look right to you?

OpenStudy (anonymous):

okay so y=-5x/9-2

OpenStudy (anonymous):

Yep, so m = -5/9. So our vector is <1, -5/9> or <9, -5>. They didn't specify a magnitude so as long as we remain proportioned, our direction is still correct.

OpenStudy (anonymous):

hmm , it says we got the wrong answer, when I put in <1,-5/9> and <9,-5>

OpenStudy (anonymous):

Hmmmmm.

OpenStudy (anonymous):

I wonder if it wants a specific magnitude...

OpenStudy (anonymous):

Hmmm actually

OpenStudy (anonymous):

I think we just take the cross product of the normal vector

OpenStudy (anonymous):

when I think about it, there is no way our answer is only a two dimensional vector.

OpenStudy (anonymous):

Yeah, I was thinking that too.

OpenStudy (anonymous):

also our line of intersection was 2D when it should be 3D...

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

So do you know how to get the normal vectors? Just get them and do the cross product.

OpenStudy (anonymous):

I can for x+y+4z=2 its <1,1,4>

OpenStudy (anonymous):

For the other? Just use distributive property to get rid of the parenthesis.

OpenStudy (anonymous):

z-x-2y=5?

OpenStudy (anonymous):

Yeah, now what is the normal vector?

OpenStudy (anonymous):

<-1,-2,1>

OpenStudy (anonymous):

So take the cross product. That is the answer.

OpenStudy (anonymous):

okay I got (9,-5,-1)

OpenStudy (anonymous):

Sweet right answer!... I hate to ask because this question was such a pain, but do you think you could help me with part b) of the question

OpenStudy (anonymous):

Ummm really quick

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

Find the equation of a plane through the origin which is perpendicular to the line of intersection of these two planes. This plane is

OpenStudy (anonymous):

This isn't any different than the problem we did in the other question. Just find the normal vectors of each plane, get the cross product. That will get you the coefficients.

OpenStudy (anonymous):

The point we know is in the plane is (0, 0, 0), since they said it passes through the origin.

OpenStudy (anonymous):

Gotta go... good luck.

OpenStudy (anonymous):

Thanks man thats great!